OpenStudy (anonymous):

One year there was a total of 84 commercial and noncommercial launches. The number of noncommercial orbital launches was 4 more than 3 times the number of commercial orbital launches. Determine the number of commercial and noncommercial launches.

OpenStudy (anonymous):

Let c=commercial and n=noncommercial. n=3c+4 c+n=84 Already have n solved, so substitute that value is second equation. c+(3c+4)=84 4c+4=84 4c=80 c=20 Solve for n using second equation. 20+n=84 n=64 There was 20 commercial launches and 64 noncommercial launches

OpenStudy (anonymous):

x = commercial, non commercial launches = 3x + 4 so x + 3x + 4 = 84 4x = 80 x = 20 so 20 commercial launches and 84-20 = 64 non commercial launches